The shareholders of Flannery Company have voted in favor of a buyout offer from Stultz Corporation. Information about each firm is given here:

Flannery Stultz
  Price–earnings ratio 13.2 37
  Shares outstanding 99,000 370,000
  Earnings $ 230,000 $ 920,000

Flannery’s shareholders will receive one share of Stultz stock for every three shares they hold in Flannery.

a-1.

What will the EPS of Stultz be after the merger? (Do not round intermediate calculations and round your answer to 3 decimal places, e.g., 32.161.)

  EPS $   
a-2.

What will the PE ratio be if the NPV of the acquisition is zero? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  PE ratio   

  

b.

What must Stultz feel is the value of the synergy between these two firms? (Do not round intermediate calculations.)
Answer #1

(a) Flannery Corporation:

PE Ratio = 13.2, Earnings = $ 230000, Shares Outstanding = 99000, EPS = 230000 / 99000 ~ $ 2.32 and Price per Share (PPS) = EPS x PE Ratio ~ $ 30.67

Stultz

PE Ratio = 37, Earnings = $ 920000, Shares Outstanding = 370000, EPS = 920000 / 370000 ~ $ 2.486 and Price per Share (PPS) = EPS x PE Ratio ~ $ 91.98 ~ $ 92  

As Flannery shareholders get one share of Stultz for three shares of their firm, Stultz needs to issue 33000 shares of its own to buyout the 99000 outstanding shares of Flannery.

Total Number of Stultz Shares post acquisition = 370000 + 33000 = 403000

Total Earnings = 230000 + 920000 = $ 1150000

EPS = 1150000 / 403000 ~ $ 2.854

(b) As Flannery shareholders receive 1 share in lieu of their three shares, Stultz actually pays ~ $ 92 for (30.67 x 3) = $ 92.01 ~ $ 92, threby establishing that the NPV of the merger (and the synergy gains) are negligible and can be assumed to be zero.

Therefore, Total Firm Value = 30.67 x 99000 + 92 x 370000 ~ $ 37076330

Total Number of Shares Outstanding = 33000 + 370000 = 403000

Price Per Share = PPS = 37076330 / 403000 ~ $ 92

PE Ratio = 92 / 2.854 ~ 32.24

(c) As NPV of the acquisition is negligible, the syngery gains can be assumed to be zero by Stultz